if f(a)=2,f'(a)=1,g(a)=-1,g'(a)=-2


then lim g(x)f(a)-g(a)f(x)/x-a as x->a?

Grade:

2 Answers

Chetan Mandayam Nayakar
12 years ago

this is an indeterminate form (0/0), thus using L'Hopital's rule lim g(x)f(a)-g(a)f(x)/x-a  as  x→a

=(f(a)g'(a)-g(a)f'(a))/1 = (-4+1)=-3 

aakanksha uday
12 years ago

Hi Ankit

solve it by dl hospital

diff. it and put the limit

you will get ans.